# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

### v1

@@ Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this by estimating the Average Treatment Effect (ATE) from decentralized observational data using federated learning, which enables inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores in a (non-)parametric manner by computing a federated weighted average of local scores, using two theoretically grounded weighting schemes—Membership Weights (MW) and Density Ratio Weights (DW)—that balance communication efficiency and model flexibility. These federated scores are then used to construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions, with theoretical analysis and experiments on simulated and real-world data highlighting their strengths and limitations relative to meta-analysis and related methods.
@@ 1 Introduction
In Randomized Clinical Trials (RCTs), treatment assignment is randomized, which ensures that the observed association between treatment and outcome reflects a causal effect. Under this design, the Average Treatment Effect (ATE) can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman et al.,, 1990), which can be further refined through covariate adjustment to reduce variance (U.S. Food and Drug Administration,, 2023; European Medicines Agency,, 2024; Lei and Ding,, 2021). However, RCTs are often expensive, time-consuming, or infeasible in practice. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán,, 2018; Hernán and Robins,, 2006). Although real-world data is abundant, drawing causal inferences from it remains challenging due to the presence of confounding covariates. As a result, the unadjusted DM estimator is biased when estimating the ATE (Grimes and Schulz,, 2002). To mitigate this bias, it is essential to adjust for confounding variables (VanderWeele,, 2019). This can be done by predicting the counterfactual outcomes of the individuals under the opposite treatment before averaging the differences between the observed and counterfactual outcomes, leading to the so-called G-Formula plug-in estimator (Robins,, 1986). Another popular method for deconfounding is to weight individuals according to their probability of receiving the treatment, thereby emulating a randomized experiment. The Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin,, 1983) relies on the estimation of the propensity score—the probability of treatment given the covariates. Doubly robust estimators, such as the Augmented IPW (AIPW) (Bang and Robins,, 2005), enhance IPW by incorporating an outcome model, offering robustness against misspecification of either model.
In causal inference, larger datasets enhance the precision and reliability of treatment effect estimates, particularly for underrepresented subgroups often overlooked in smaller studies. However, real-world data is typically decentralized—collected across several hospitals, companies, or countries—making it challenging to aggregate into a centralized dataset. This difficulty is particularly acute in domains like healthcare, where privacy concerns, regulatory barriers, data ownership and logistical issues (such as responsibility for data storage and governance) complicate data sharing. Federated Learning (FL) (Kairouz et al.,, 2021) offers a promising solution by enabling model training directly on decentralized data within a server-client architecture, without sharing individual-level data. By exchanging only aggregate statistics such as model updates or gradients, FL has the potential to support collaborative causal inference studies with performance approaching that of centralized analyses. However, realizing this potential requires tailored methodological developments.
Contributions. In this article, we build on ideas from Federated Learning to design (A)IPW estimators for decentralized observational data, moving beyond the simple aggregation of local ATE estimates used in traditional meta-analysis approaches.
@@
At the core of our approach is a flexible, nonparametric strategy for federating propensity scores. Unlike prior methods that aggregate parameters of a global parametric model via “one-shot” averaging (Xiong et al.,, 2023) or federated gradient descent (Khellaf et al.,, 2025), we construct a global propensity score model as a mixture of locally estimated models. This is done through a two-step process: (1) local estimation of propensity scores at each site, which naturally accommodates site-specific heterogeneity in treatment assignment mechanisms and offers flexibility in model choice, and (2) aggregation of these local models into a global one using carefully selected federated weights. We introduce two theoretically grounded weighting schemes: Membership Weights (MW), which represent the probability of an individual belonging to each site given their covariates, and Density Ratio Weights (DW), which model the relative density of covariates at each site compared to the global population. We explain how to infer MW and DW weights in a federated manner and highlight the distinct trade-offs they offer between communication efficiency and modeling assumptions. Using these federated propensity scores, we construct two ATE estimators: the Federated Inverse Propensity Weighting estimator (Fed-IPW) and its augmented variant (Fed-AIPW). We derive asymptotic variance expressions for these estimators and compare them to the meta-analysis estimators—often overlooked in FL (Khellaf et al.,, 2025)––, proving that our estimators achieve lower or equal variances in all scenarios considered.
@@
Our approach is particularly beneficial in scenarios with poor or nonexistent overlap between treated and control groups within individual sites. In such cases, cross-site collaboration becomes crucial, as propensity score functions vary across sites. Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup that remains untreated elsewhere—the global dataset achieves markedly greater overlap, providing essential support for estimating treatment effects that would otherwise be poorly estimated within isolated sites. Additionally, our framework naturally accommodates a wide range of cross-site heterogeneities, including variations in sample sizes, treatment policies, covariate distributions, and violations of positivity. Numerical experiments on both simulated and real-world data validate our theoretical findings and highlight the practical value of our method.
@@ 2.1 ATE estimators from centralized multi-site observational data
In this section, we recall the key components of ATE estimation in a centralized multi-site setting. Following the potential outcomes framework (Rubin,, 1974; Splawa-Neyman et al.,, 1990), we consider random variables $(X,H,W,Y{(1)},Y{(0)})$, where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively. We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is $Y=WY(1)+(1-W)Y(0)$, and that potential outcomes are uniformly bounded.
In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$. We aim to estimate the ATE defined as the risk difference $\tau=\mathbb{E}\left[Y(1)-Y(0)\right]=\mathbb{E}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$, where the expectation is taken over the population $\mathcal{P}$. To be able to identify the ATE, we assume unconfoundedness.
@@
While Assumption 1 is a standard assumption in causal inference, we further consider 2, which is specific to the multi-site setting. Together with Assumption 1, it ensures that the variables $X$ form a sufficient set of covariates to adjust for confounding.
@@
Our setting is depicted in the graphical model in Figure 1, which highlights that we eliminate any direct effect of the site on the outcome.
$W$$X$$Y$$H$
@@
We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$. The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager,, 2024).
Assumption 3 (Global overlap).
$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\big{]}<\infty$.
Assumption 3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments. A lower value of $\mathcal{O}_{\mathrm{global}}$ indicates that these probabilities lie further away from 0 and 1, which corresponds to better overlap. For further insights on overlap, see Li et al., 2018a ; Li et al., 2018b , and for a “misoverlap” metric, refer to Clivio et al., (2024).
With Assumptions 1, 2 and 3, the ATE is identifiable as $\tau=\mathbb{E}\big{[}\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\big{]}$ (see Appendix A.1). We denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript . We define the centralized, multi-site estimators as follows.
Definition 1 (Oracle multi-site centralized estimators).
@@
where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$.
These oracle estimators are unbiased and asymptotically normal.
Theorem 1 (Consistency of oracle multi-site centralized estimators).
Under Assumptions 1, 2 and 3, we have the following asymptotic normality results:
$\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V^{\star})\ \text{with}$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{*}\right]=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},$
$\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{*}\right]=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right],$
where $\tau(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ denotes the Conditional Average Treatment Effect (CATE).
These asymptotic variances align with those in the single-site setting (Hirano et al.,, 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the AIPW estimator is doubly robust, meaning it remains consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al.,, 2018).
@@ 2.2 Meta-analysis estimators
We now turn to a decentralized setting in which the $K$ sites cannot share individual-level data. A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al.,, 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation. We will need the following assumption.
@@
$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big{[}(e(X)(1-e(X)))^{-1}\mid H=k\big{]}<\infty$.
Assumption 4 is much stronger than its global counterpart (Assumption 3), as it requires the overlap to hold locally for each site. Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows.
Definition 2 (Oracle meta-analysis estimators).
@@
where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$.
While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau$, which decomposes as $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}\left[Y(1)-Y(0)\mid H=k\right]$ denotes the local ATE at site $k$. This bias arises whenever the $\{\tau_{k}\}_{k}$ differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates).
Theorem 2 (Consistency of oracle meta-analysis estimators).
@@
$\mathbb{V}\left[\hat{\tau}_{\mathrm{IPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right],\quad\quad\mathbb{V}\left[\hat{\tau}_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}\right]=\frac{1}{n}\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\frac{1}{n}\mathbb{V}\left[\tau_{H}\right]$
@@
with within-site variances $V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}$, $V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right]$ and between-sites variance of the local ATEs $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$.
@@
This result is proved in Appendix A.3. A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated in practice—for example, when a site follows a deterministic treatment policy (e.g., treating all patients or only those with a specific condition). In such cases, these estimators become ill-defined and yield biased estimates of the global ATE.
To overcome this, we introduce a federated approach that estimates the global propensity score $e$ as a weighted combination of the local scores $e_{k}$, enabling valid inference even without local overlap.
3 Related work
Federated causal inference has only recently gained attention. Khellaf et al., (2025) propose a federated framework for estimating the ATE across multiple RCTs by fitting separate parametric linear outcome models for treated and control groups at each site using federated learning, followed by ATE estimation via the G-formula. Their method is motivated by the fact that, in RCTs, this estimator remains unbiased under model misspecification and achieves lower variance than the Difference-in-Means estimator.
In observational studies, however, outcome model misspecification can bias ATE estimates. To address this, Vo et al., (2022) introduce a Bayesian framework using Gaussian processes trained via federated gradient descent. Their method requires estimating a multi-site covariance kernel, which involves each site computing and sharing the first four moments of its data—a computationally intensive process that raises concerns regarding scalability and communication efficiency.
Alternative approaches, such as IPW, require estimating the propensity score. Guo et al., (2025) propose a federated method for learning a global propensity score model using a consensus voting scheme over parametric model parameters, retaining only sites that conform to a shared specification. In contrast, our approach does not assume a common propensity score function: each site can adopt its own, potentially distinct model.
To address site-level heterogeneity, Xiong et al., (2023) assume a parametric logistic propensity score model with shared and site-specific parameters, federating only the common ones in a single round of communication round. In a similar line of work, Yin et al., (2025) estimate a global parametric propensity score model that adjusts for both covariates and site membership, but their handling of heterogeneity is limited to a site-specific scalar, restricting the model’s flexibility. In contrast, our method makes no structural assumptions about the form of the propensity score. We enable fully nonparametric global estimation while allowing flexible and heterogeneous local models, without requiring prior knowledge of shared or site-specific components. Crucially, unlike the above methods, which assume overlap holds locally at each site, our approach explicitly relaxes this strong assumption.
Although not always framed within a federated context, a related body of work focuses on the generalization of causal findings from source data to a target population. Assuming homogeneous outcome models and propensity scores across sites, Han et al., (2025) propose a density ratio weighting method to generalize ATE estimates from multiple source sites to a target domain. Their approach resembles meta-analysis, relying on aggregate statistics—such as sample sizes and local ATE variances—rather than individual-level data. While they estimate density ratios using parametric models, our Membership Weights approach leverages supervised learning methods (e.g., logistic regression) that have lower sample complexity and make less stringent assumptions on the sites’ distributions. Similarly, Guo et al., (2024) address generalization to a target population composed of multiple source sites and previously excluded individuals. Their method assumes the target propensity score is proportional to those of the sources and estimates local scores via density ratios between treated or control subgroups and the target population. However, nonparametric estimation of these ratios requires sufficient data in each treatment arm at every site—a condition often unmet in small or imbalanced settings. Furthermore, the authors do not address how this could be implemented in a federated setting.

### v2

@@ Abstract
Causal inference typically assumes centralized access to individual-level data. Yet, in practice, data are often decentralized across multiple sites, making centralization infeasible due to privacy, logistical, or legal constraints. We address this problem by estimating the Average Treatment Effect (ATE) from decentralized observational data via a Federated Learning (FL) approach, allowing inference through the exchange of aggregate statistics rather than individual-level data. We propose a novel method to estimate propensity scores by computing a federated weighted average of local scores with Membership Weights (MW)—probabilities of site membership conditional on covariates—which can be flexibly estimated using parametric or non-parametric classification models. Unlike density ratio weights (DW) from the transportability and generalization literature, which either rely on strong modeling assumptions or cannot be implemented in FL, MW can be estimated using standard FL algorithms and are more robust, as they support flexible, non-parametric models—making them the preferred choice in multi-site settings with strict data-sharing constraints. The resulting propensity scores are used to construct Federated Inverse Propensity Weighting (Fed-IPW) and Augmented IPW (Fed-AIPW) estimators. Unlike meta-analysis methods, which fail when any site violates positivity, our approach leverages heterogeneity in treatment assignment across sites to improve overlap. We show that Fed-IPW and Fed-AIPW perform well under site-level heterogeneity in sample sizes, treatment mechanisms, and covariate distributions. Both theoretical analysis and experiments on simulated and real-world data highlight their advantages over meta-analysis and related methods.
@@ 1 Introduction
The Average Treatment Effect (ATE) is a key causal estimand used to quantify the effect of a treatment on an outcome and is commonly employed as the primary measure of efficacy in evaluating new therapies, including vaccines, before regulatory approval (Polack et al., 2020). In Randomized Clinical Trials (RCTs), treatment assignment is randomized, ensuring that the observed association between treatment and outcome reflects a causal effect. Under this design, the ATE can be consistently estimated using a simple Difference-in-Means (DM) estimator (Splawa-Neyman, 1990), which can be further refined through covariate adjustment to reduce variance (FDA, 2023; EMA, 2024; Lei and Ding, 2021). However, RCTs are often expensive, time-consuming, or infeasible. In such cases, estimating treatment effects from observational data becomes the only viable alternative (Hernán, 2018; Hernán and Robins, 2006). Although such real-world data is abundant, drawing causal inferences from it is challenging due to confounding covariates, rendering the unadjusted DM estimator biased (Grimes and Schulz, 2002). Adjusting for confounders is thus essential (VanderWeele, 2019). This can be done by predicting counterfactual outcomes before averaging the differences (the G-formula plug-in estimator, Robins, 1986). Another approach is to weight individuals according to their treatment probability, emulating a randomized trial. For instance, the Inverse Propensity Weighting (IPW) estimator (Rosenbaum and Rubin, 1983) relies on estimating the propensity score—the probability of treatment given covariates. Doubly robust estimators such as the Augmented IPW (AIPW) (Bang and Robins, 2005) combine weighting with outcome modeling to remain consistent as long as either model is correctly specified.
Larger datasets improve the precision of treatment effect estimates, especially for underrepresented subgroups. Yet real-world data is typically decentralized—spread across hospitals, companies, or countries—making aggregation difficult, particularly in healthcare where privacy regulations, data ownership, and governance issues impede centralization. Federated Learning (FL) (Kairouz et al., 2021) offers a solution to train models across distributed data without sharing individual-level data. While FL has been largely applied to prediction tasks, its extension to causal inference remains limited. This problem is especially challenging in observational studies, where differences in covariate distributions and treatment assignment mechanisms across sites create multiple sources of heterogeneity that must be addressed without sharing raw data, while achieving results comparable to centralized analyses—an issue that remains largely unsolved.
@@
Contributions. We propose federated (A)IPW estimators for decentralized observational data, moving beyond the aggregation of local ATE estimates used in meta-analysis (Riley et al., 2023). At the core of our approach is a flexible, primarily non-parametric strategy for federating propensity scores. Unlike prior methods that fit a single global parametric model via parameter averaging (Xiong et al., 2023) or federated gradient descent (Guo et al., 2025), we construct a global propensity score as a mixture of locally estimated models. This involves (1) local estimation of propensity scores at each site—accommodating heterogeneity in treatment assignment and flexibility in model choice—and (2) aggregation into a global score using Membership Weights (MW), i.e., the probability of site membership given covariates. MW can be estimated in a federated manner using flexible, potentially non-parametric classification models, ensuring both robustness and communication efficiency. In contrast, transferring Density Ratio Weights (DW) from the transportability and generalization literature to our setting requires strong modeling assumptions, as they are otherwise incompatible with FL constraints. Using our federated propensity scores, we build the Federated IPW (Fed-IPW) and its augmented variant (Fed-AIPW), derive their variances, and show they achieve equal or lower variance than meta-analysis estimators.
@@
Our approach is particularly advantageous when overlap between treatment groups is poor or absent within sites. In such scenarios, cross-site collaboration becomes crucial, as combining sites increases overall overlap and enables treatment effect estimation that may be infeasible locally. Indeed, when treatment assignment mechanisms differ substantially—such as when one site treats a subgroup absent elsewhere—the combined dataset achieves markedly greater overlap, allowing treatment effects to be estimated that would otherwise be poorly identified in isolation. Additionally, our framework naturally accommodates heterogeneity in sample sizes, treatment policies, covariate distributions, and violations of positivity. Numerical experiments on simulated and real data confirm our theoretical findings and highlight the method’s practical benefits.
Related work. Federated causal inference is still nascent. Khellaf et al. (2025) estimate federated G-formula ATE estimates across multiple RCTs by fitting parametric outcome models at each site via FL, achieving lower variance than DM. In observational settings, Vo et al. (2022) propose a federated Bayesian approach using Gaussian processes with a shared covariance kernel, but it requires sharing the first four moments of the data, limiting scalability and efficiency. Guo et al. (2025) learn a global propensity score via consensus voting over parametric parameters, retaining only sites meeting a shared specification. In contrast, our method assumes no common propensity score: each site may fit its own model.
To address site heterogeneity, Xiong et al. (2023) use a logistic propensity model with shared and site-specific parameters, federating only the common ones. Yin et al. (2025) fit a global model adjusting for covariates and site membership but limit heterogeneity to a site-specific scalar. By contrast, our method makes no structural assumptions, enabling fully nonparametric estimation with heterogeneous local models and relaxing the need for local overlap at each site.
A related body of work focuses on generalizing causal findings from multi-site source populations to a target population. Han et al. (2025) use density ratio weighting of local ATEs to adjust for covariate shift but assume homogeneous nuisance functions across sites and rely on meta-analysis of aggregate statistics. Guo et al. (2024) extend this idea by applying density ratio weights to aggregate local propensity scores to construct a target-specific score, which requires density estimation within each treatment arm at each site—demanding large sample sizes per arm for stable estimates. In both cases, non-parametric density ratio estimation is infeasible under FL constraints, as it requires sharing raw data or detailed covariate representations (e.g., kernel evaluations or high-dimensional histograms). In contrast, our MW-based approach leverages flexible parametric or non-parametric supervised models (e.g., logistic regression, neural networks, gradient-boosted trees) for which federated training is already established, operating without sharing raw data, potentially at lower sample complexity, with full support for heterogeneous nuisance functions, and allowing the natural inclusion of external control arms.
@@ 2.1 ATE Estimators from Centralized Multi-Site Observational Data
In this section, we recall the key components of ATE estimation in a centralized multi-site setting. Following the potential outcomes framework (Rubin, 1974; Splawa-Neyman, 1990), we consider random variables (X,H,W,Y(1),Y(0))(X,H,W,Y{(1)},Y{(0)}), where $X\in\mathbb{R}^{d}$ represents patient covariates, $H\in[K]$ indicates site membership, $W\in\{0,1\}$ denotes the binary treatment, and $Y(1)$ and $Y(0)$ are the potential outcomes under treatment and control, respectively. We assume that the Stable Unit Treatment Values Assumption (SUTVA) holds, so that the observed outcome is Y=WY(1)+(1W)Y(0)Y=WY(1)+(1-W)Y(0), and that the potential outcomes are uniformly bounded. In the centralized setting, we observe $n=\sum_{k=1}^{K}n_{k}$ observations of independently and identically distributed (i.i.d.) tuples $(H_{i},X_{i},W_{i},Y_{i})_{i\in 1,\dots,n}\sim\mathcal{P}^{\otimes n}$, with $n_{k}=\sum_{i=1}^{n}\mathds{1}_{\{H_{i}=k\}}$ the number of observations in site $k$. We aim to estimate the ATE defined as the risk difference $\tau=\mathbb{E}\left[Y(1)-Y(0)\right]=\mathbb{E}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$, where the expectation is taken over the population $\mathcal{P}$. To be able to identify the ATE, we assume unconfoundedness (standard in causal inference) and further consider ˜2, which is specific to the multi-site setting.
@@
Robertson et al. (2021) refer to ˜2 as the no center-outcome association condition. It can be stated as a testable null hypothesis requiring that, for every treatment level $w$ and any pair of centers $(k,k^{\prime})$, $\mathbb{E}[Y\mid X,W=w,H=k]=\mathbb{E}[Y\mid X,W=w,H=k^{\prime}]$. Combined with Assumption 1, this ensures that $X$ forms a sufficient set of covariates for confounding adjustment. Our setting is depicted in the graphical model in Figure 1, highlighting that we remove any direct effect of the site on the outcome.
We define $\mu_{w}(x)=\mathbb{E}\left[Y\mid X=x,W=w\right]$ for $w\in\{0,1\}$, and let $\tau(x)=\mu_{1}(x)-\mu_{0}(x)=\mathbb{E}[Y(1)-Y(0)\mid X=x]$ be the Conditional Average Treatment Effect (CATE). The oracle propensity score is denoted by $e(x)=\mathbb{P}(W\mid X=x)$, and we consider the weak (global) overlap assumption (Wager, 2024).
Assumption 3 (Global overlap).
$\mathcal{O}_{\mathrm{global}}=\mathbb{E}\big[(e(X)(1-e(X)))^{-1}\big]<+\infty$.
WWXXYYHH
@@
˜3 is crucial for propensity score-based estimators, as it states that every region of the covariate space has a non-zero probability of receiving both treatments. A lower value of $\mathcal{O}_{\mathrm{global}}$ indicates that these probabilities lie further away from 0 and $1$, which corresponds to better overlap. For further insights on overlap, see Li et al. (2018a, b), and for a “misoverlap” metric, refer to Clivio et al. (2024).
With Assumptions 1, 2 and 3, the ATE is identifiable as $\tau=\mathbb{E}\big[\frac{WY}{e(X)}-\frac{(1-W)Y}{1-e(X)}\big]$ (see Appendix A.1). Throughout the paper, we denote oracle ATE estimators, which assume knowledge of the nuisance components $e,\mu_{0},\mu_{1}$, by a superscript . We define the Oracle multi-site centralized estimators as follows:
@@
where $\tau_{\mathrm{IPW}}(X_{i};e)=\frac{W_{i}Y_{i}}{e(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-e(X_{i})}$ and $\tau_{\mathrm{AIPW}}(X_{i};e,\mu_{1},\mu_{0})=\mu_{1}(X_{i})-\mu_{0}(X_{i})+\frac{W_{i}(Y_{i}-\mu_{1}(X_{i}))}{e(X_{i})}-\frac{(1-W_{i})(Y_{i}-\mu_{0}(X_{i}))}{1-e(X_{i})}$. These oracle estimators are unbiased and asymptotically normal.
Theorem 1.
Under Assumptions 1, 2 and 3, we have $\sqrt{n}(\hat{\tau}-\tau)\to\mathcal{N}(0,V)$ with
$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}=\mathbb{E}\left[\frac{Y(1)^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e(X)}\right]-\tau^{2},\\ V_{\mathrm{AIPW}}=\mathbb{V}\left[\tau(X)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e(X)}\right)\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e(X)}\right)^{2}\right].\end{array}\right$.
The above asymptotic variances align with those in the single-site setting (Hirano et al., 2003), as detailed in Appendix A.2. However, in practice, the propensity score and outcome models are typically unknown and must be estimated from data. This creates a challenge in the decentralized setting, where centralizing data to compute $\mu_{1}$, $\mu_{0}$, and $e$ is not feasible. Therefore, the estimators in Definition 1 need to be adapted to this setting. Importantly, the (non-oracle) AIPW estimator is inherently doubly robust, remaining consistent as long as either the outcome or the propensity score model is correctly specified (Chernozhukov et al., 2018).
@@ 2.2 Meta-Analysis Estimators
We now turn to a decentralized setting in which the $K$ sites cannot share individual-level data. A natural baseline for estimating the ATE across sites is a two-stage meta-analysis approach (Burke et al., 2017), wherein each site independently estimates the relevant nuisance parameters and communicates only the resulting ATE estimates for aggregation. In this setting, we need the following assumption.
@@
$\forall k\in[K],\mathcal{O}_{k}=\mathbb{E}\big[(e(X)(1-e(X)))^{-1}\mid H=k\big]<+\infty$.
Assumption 4 is much stronger than global overlap (˜3), as it must hold at every site. Denoting by $e_{k}(x)=\mathbb{P}(W=1\mid X=x,H=k)$ the oracle local propensity score at site $k$, we can define the oracle meta-analysis estimators as follows:
@@
where $\hat{\tau}_{\mathrm{IPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{IPW}}(X_{i};e_{k})$ and $\hat{\tau}_{\mathrm{AIPW}}^{(k)}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\tau_{\mathrm{AIPW}}(X_{i};e_{k},\mu_{1},\mu_{0})$ are the local estimators at site $k$. While alternative aggregation weights—such as the inverse variance of local estimates—can be considered, they produce, in our setting, biased estimates of the global ATE $\tau=\sum_{k=1}^{K}\rho_{k}\tau_{k}$, where $\rho_{k}=\mathbb{P}(H=k)$ and $\tau_{k}=\mathbb{E}[Y(1)-Y(0)\mid H=k]$ is the local ATE. This bias appears whenever the $\tau_{k}$} differ, which commonly occurs when covariate distributions vary across sites and treatment effects are heterogeneous (i.e., depend on covariates), see (Berenfeld et al., 2025).
Theorem 2.
@@
$V_{\mathrm{IPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{IPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],\quad\quad V_{\mathrm{AIPW}}^{\mathrm{meta^{*}}}=\sum_{k=1}^{K}\rho_{k}V_{\mathrm{AIPW}}^{(k)}+\mathbb{V}\left[\tau_{H}\right],$
with within-site variance
@@
$\left\{\begin{array}[]{l}V_{\mathrm{IPW}}^{(k)}\!=\mathbb{E}\left[\frac{Y(1)^{2}}{e_{k}(X)}\mid H=k\right]\!+\mathbb{E}\left[\frac{Y(0)^{2}}{1-e_{k}(X)}\mid H=k\right]\!-\tau_{k}^{2}\\ V_{\mathrm{AIPW}}^{(k)}=\mathbb{V}\left[\tau(X)\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{1}(X))^{2}}{e_{k}(X)}\right)^{2}\mid H=k\right]+\mathbb{E}\left[\left(\frac{(Y-\mu_{0}(X))^{2}}{1-e_{k}(X)}\right)\mid H=k\right],\end{array}\right$.
@@
and $\mathbb{V}\left[\tau_{H}\right]=\mathbb{V}\left[\mathbb{E}\left[Y(1)-Y(0)\mid H\right]\right]$ the between-sites variance of the local ATEs.
@@
This result is proved in Appendix A.3. A key limitation of meta-analysis estimators is their reliance on Assumption 4, which is fragile and often violated—for instance, when a site applies a deterministic treatment policy (treating all patients or only a subgroup). In such cases, these estimators are ill-defined, yielding biased ATE estimates. To address this, we propose a federated approach that constructs the global propensity score $e$ as a weighted combination of local scores $e_{k}$, enabling valid inference even without local overlap.